OpenStudy (anonymous):

A circle is described by the equation x2 + y2 + 14x + 2y + 14 = 0. What are the coordinates for the center of the circle and the length of the radius? A)(-7, -1), 36 units B)(7, 1), 36 units C)(7, 1), 6 units D)(-7, -1), 6 units help me plz

OpenStudy (bradely):

x^2 + y^2 + 14x + 2y + 14 = 0. is compare with x^2 + y^2 + 2gx + 2fy + c= 0. Center (-g,-f) Radius r =sqrt(g^2+f^2-c) 2g =14 g=7 2f= 2 f =1 c=14 Center, C =(-g,-f) =(-7,-1) Radius r= sqrt(7^2+1^2-14) =sqrt(49+1-14) =sqrt(36) =6 Option D is correct Source: http://www.mathskey.com/question2answer/